OpenStudy (anonymous):

Simplify the quotient (4.5 x 〖10〗^3)/(9 x 〖10〗^7 ). Write your answer in scientific notation. Show your work

OpenStudy (anonymous):

im about to

OpenStudy (anonymous):

k thx

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

thx now im gonna answer

OpenStudy (anonymous):

4.5×1039×107 =0.5*10^3-7 =0.5*10^-4 =5*10^-1*10^-4 =5*10^-1-4 =?

OpenStudy (anonymous):

2495 in scientific notation u should know how to do that

OpenStudy (anonymous):

medal plz thx in advance

OpenStudy (anonymous):

@jigglypuff314 please do something @clanger555 doesnt stop doing this

OpenStudy (anonymous):

@jigglypuff314 Clanger is giving direct answers and i will tag u in my question, they were spamming my question with the answer multiple times and it gets annoying